OpenStudy (kidthatbro8):

Winnie poured 14 cups of water into a rectangular container measuring 13 inches by 7 inches by 6 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points)

OpenStudy (kidthatbro8):

please, help me.

OpenStudy (igreen):

Volume of a rectangular Prism: \(V = LWH\)

OpenStudy (igreen):

Change 6 centimeters to inches: \(6cm = 2.36 in\) \(V = (13)(7)(2.36)\) Multiply from left to right: \(V = (91)(2.36)\) \(V = 214.76\) So the volume of the rectangular prism is 214.76.

OpenStudy (igreen):

For Part B: 1 cup = 14.44 cubic inches. 14 cups total. So: 14 * 14.44 = 202.16. So that's 202.16 cubic inches.

OpenStudy (igreen):

For Part C: \(V = LWH\) Plug in what we know: \(202.16 = (13)(7)h\) Solve for h: Multiply: \(202.16 = 91h\) Divide 91 to both sides: \(h \approx 2.22\)

OpenStudy (kidthatbro8):

thank you.

OpenStudy (igreen):

No problem.